acidic ingredient in vinegar is acetic acid. The pH of vinegar is around 2.4, and the 

molar concentration of acetic acid in vinegar is around 0.85 M. Based on this 

information, determine the value of the acid ionization constant, Ka, for acetic acid

a) 2.5 × 10–5

b) 5.0 × 10–5

c) 4.7 × 10–3

d) None of the above


1
Expert's answer
2020-07-14T15:17:03-0400

H2O(l) + HAcO(aq) "\iff" H3O+(aq) + AcO(aq)

I 0.85M 0M 0M

C -xM +xM +xM

E (0.85-x) xM xM


ka = x2/ (0.85-x)

x = [H3O+] = 10-pH-

x = 10-2.4

x = 0.00398

ka = (0.00398)2/0.85

ka = 1.9 x 10-5

The value of the acid ionization constant, Ka, for acetic acid is 1.9 x 10-5